What is the density of a piece of aluminium with a volume of 5.60 x 10-6m3 and a mass of 15.1 g?
1
Expert's answer
2020-10-01T10:07:52-0400

We can find the density of a piece of aluminium from the formula:


"\rho = \dfrac{m}{V} =\dfrac{15.1 \ g \cdot (\dfrac{1 \ kg}{1000 \ g})}{5.60 \cdot 10^{-6} \ m^3} = 2696 \ \dfrac{kg}{m^3}."

Answer:

"\rho = 2696 \ \dfrac{kg}{m^3}."